OpenStudy (anonymous):

Find the volume: Under z =x^2+y^2 and above the region R={(x,y)| x^2+ y^2≤ 4} Please explain step by step.

OpenStudy (experimentx):

the graph is going to be 3d-parabolic we have to find the vertex first.

OpenStudy (experimentx):

looks like 0,0,0 is going to be vertex

OpenStudy (anonymous):

follow this !! www.physicsforums.com › ... › Calculus & Beyond

OpenStudy (anonymous):

See a couple of solved example from my site http;//www.saab.org You can practice similar problems on http://moltest.missouri.edu/mucgi-bin/calculus.cgi Do not forget to like http;//www.saab.org on Facebook

OpenStudy (turingtest):

ok, right off the bat I'm thinking cylindrical coordinates

OpenStudy (turingtest):

@Sillem do you know how to convert this into cylindrical coordinates?

OpenStudy (turingtest):

If you convert this to cylindrical coordinates you will find that the region is bound by\[0\le\theta\le2\pi\]\[0\le r\le2\]and converting the equation we have for the parabaloid we get the bounds\[0\le z\le r^2\]remembering that \(dV=rdrd\theta dz\) in cylindrical coordinates we now just have to integrate\[\int\int\int dV\]along the prescribed bounds, and in an order that makes some kind of sense. I'll leave that final task to you.

OpenStudy (turingtest):

having a problem? please tell me where